OpenStudy (anonymous):

NOW LAST QUESTION The velocity of sound in air is given by the equation______ , where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the temperature when the velocity is 329 meters per second by graphing the equation. Round the answer to the nearest degree.

OpenStudy (anonymous):

\[v = 20 \sqrt{273 + t}\]

OpenStudy (anonymous):

This looks like normal substitution... Is it as straightforward as 20 * sqrt (273 + 329) ?

OpenStudy (anonymous):

oops, sorry, read that problem wrong. that's incorrect ^^

OpenStudy (anonymous):

329 = 20*sqrt(273 + t)

OpenStudy (anonymous):

what is sqrt?

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

o ok :)

OpenStudy (anonymous):

so, you just want to solve for t: (draw the square root symbol instead of writing "sqrt") 329 = 20*sqrt(273 + t) Divide both sides by 20 16.45 = sqrt(273 + t) Then you can square both sides to get rid of the square root. 16.45^2 = 273 + t

OpenStudy (anonymous):

So punch 16.45 into your calculator, multiply it by 16.45, and then subtract 273 to solve for the temperature t.

OpenStudy (anonymous):

And then your problem says round that to the nearest whole degree.

OpenStudy (anonymous):

So it would be -6?